The charges to Work in Process-Assembly Department for a period, together with information concerning production, are as follows. All direct materials are placed in process at the beginning of production Work in Process-Assembly Department Bal., 3,000 units, 40% completed Direct materials, 71,000 units @ $1.3 Direct labor Factory overhead Bal. 2 units, 80% completed 6,300 92,300 To Finished Goods, 69,000 units 113,700 44,260 Determine the following a. The number of units in work in process inventory at the end of the period. units b. Equivalent units of production for direct materials and conversion. If an amount is zero or a blank, enter in "O" Work in Process-Assembly Department Equivalent Units of Production for Direct Materials and Conversion Costs Whole Units Equivalent Units Direct Materials Equivalent Units Conversion Inventory in process, beginning Started and completed Transferred to finished goods Inventory in process, ending
Total units c. Costs per equivalent unit for direct materials and conversion. If required, round your answers to the nearest cent. Costs Per Equivalent Unit Direct Materials Conversion d. Cost of the units started and completed during the period
Answer #1
a
Inventory in process, beginning 3000
Add: Units started 71000
Less: Transferred to finished goods -69000
Units in work in process at the end 5000 units
b
Whole units Equivalent units
Direct materials
Equivalent units
Conversion
Inventory in process, beginning 3000 0 1800
Started and completed 66000 66000 66000
Transferred to finished goods 69000 66000 67800
Inventory in process, ending 5000 5000 4000
Total units 74000 71000 71800
c
Costs per equivalent unit
Direct materials 1.30 =92300/71000
Conversion 2.20 =(113700+44260)/71800
d
Units Started and completed 66000
X Cost per equivalent unit 3.50 =1.3+2.2
Cost of units started and completed 231000